OpenStudy (anonymous):

Evaluate the integral without using a calculator: [-1(bottom number),1(top number)] dx/(1+x^2)

OpenStudy (anonymous):

why tan(u)?

OpenStudy (anonymous):

I dont remember them... what are all the trigo. subs

OpenStudy (stamp):

\[\int\limits_{-1}^{1}\frac{ dx }{ 1+x^{2} }=tan^{-1}(b) - \tan^{-1}(a)\] \[\ =tan^{-1}(1) - \tan^{-1}(-1) \] \[\frac{ \pi }{ 4 }- \frac {-\pi}{ 4 } = \frac{ \pi }{ 4 }+\frac{ \pi }{ 4 } = \frac{2 \pi}{4}\] \[=\frac{\pi}{2} \]